OpenStudy (faman39):

Match each actual value with the predicted value that results in a 15% error. a. Predicted : 119 b.Predicted : 85 c. Predicted : 102 d. Predicted :92 Actual :80 Actual :120 Actual :100 Actual : 140

OpenStudy (anonymous):

if you match any of the predicted values with any of the actual values, wouldnt they all be errors? because none are the same. I think what this problem is looking for is that you pick the actual value thats closest to the predicted values (and repeat for all 4 predictions) but otherwise im not exactly sure what the question is asking

OpenStudy (anonymous):

Is there more to this question? Such as, there supposed to being a specific error percentage, or error pattern. Anything like this?

OpenStudy (anonymous):

Can you reply?

OpenStudy (faman39):

I m sorry guys, i put wrong

OpenStudy (faman39):

before

OpenStudy (faman39):

Now i fixed the question

OpenStudy (anonymous):

in that case: a=102, b=100, c=120, and d =80. note these arent exactly 15% but theyre the closest thing considering the limited data given

OpenStudy (faman39):

Thank you guys !

hero (hero):

Percent Change Formula: \[\% \Delta = \frac{|a - b|}{a} \times 100\] a = 1st Number b = 2nd Number Try different values until you find the right combination

hero (hero):

When I tried, I got 85 Predicted, 100 Actual

OpenStudy (faman39):

Thank you so much :)